A dam holds back the water in a lake. If the dam has a small hole, 1.4 meters below the surface of the lake, then at what speed does water exit the hole?


1
Expert's answer
2021-02-28T07:32:02-0500
"v=\sqrt{2gh}\\v=\sqrt{2(9.8)(1.4)}=5.2\frac{m}{s}"